OpenStudy (anonymous):

Evaluate: 32 + (7 − 2) • 4 − 6/3

OpenStudy (anonymous):

Your answer is 50.

OpenStudy (nenadmatematika):

32+5*4-2=52-2=50

OpenStudy (anonymous):

Theirs no 50.

OpenStudy (anonymous):

24 54 27 28

OpenStudy (anonymous):

Then either your choices are wrong or you accidentally entered the wrong numbers.

OpenStudy (anonymous):

Wait.. I wrote it wrong. I'm sorry!

OpenStudy (anonymous):

haha :)